OpenStudy (anonymous):

if the function is defined by f(x)=(ax-2)/(2x-3), x is not equal to k, satisfies f(x)=x for all real values of x except k (i) State the value of k (ii) State the range of values of a (i)f(x)=(ax-2)/(2x-3) ax-2/2x-3=x x(2x-3)=ax-2 x(2x-3-a)=-2 2=x(a+3-2x)

OpenStudy (anonymous):

Well then, if i'm reading you right\[2x-3=0\rightarrow x=3/2\rightarrow k= 3/2\]And then,\[x=(ax-2)/(2x-3)\rightarrow x(2x-3)=ax-2 \rightarrow ax = 2x ^{2}-3x+2\]So... a can be equal to pretty much anything. Hope i could help, didn't really got your question. Cheers.